Matt recently deposited $27,500 in a savings account paying a guaranteed interest rate of 5.2% for the next 10 years if Matt expects his marginal rate tax to be 22% for the next 10 years how much interest will he earn after-tax for the first year of his investment? how much interest will I earn after-tax for the second year of his investment if he withdraws enough cash every year to pay the tax on the interest he earns? How much will he have in the account after 4 years? How much will he have in the account after 7 years?
Answer #1

1. First year returns => 27500*5.2%*(1-0.22) = $ 1,115.4

2. Second year => (27,500+1115.40)*5.2%*(1-0.22) = $ 1,160.64

3.

Interest after tax => (5.2%)(1-0.22) => 4.056%

Balance after 4 years: (27500)*(1+(4.056/100))^4 = $ 32,240.46

4. Account balance after 7 years: (27500)*(1+4.056/100)^7 = $ 36,324.75